# Sailor: Automating Distributed Training over Dynamic, Heterogeneous, and Geo-distributed Clusters

### v1

@@ 1. Introduction
GPUs are in high demand for large-scale Machine Learning (ML). As ML models continue to grow exponentially in size, they require an increasing number of GPUs to train and fine-tune. This high demand makes it difficult for model developers to allocate the desired number of accelerators to train models at high throughput in public clouds or enterprise clusters (Um24Metis, ; strati24crossregion, ; Zongheng23Skypilot, ). Datacenters typically host a variety of GPU types and generations, spread across geographic regions (gcpgpus, ; Jia2022Whale, ; weng22mlaas, ; Jeon19philly, ; wu2024falcon, ). Yet model developers tend to restrict model training to homogeneous clusters of GPUs, since state-of-the-art distributed training frameworks like Megatron-LM (shoeybi2020megatronlm, ) and DeepSpeed (megatron-deepspeed, ) assume homogeneous GPUs and inter-node bandwidth. The demand for large, homogeneous GPU clusters compounds the scarcity of high-end GPUs.
Allowing a training job to run on heterogeneous GPU types and/or GPUs distributed across zones (i.e., with heterogeneous inter-node bandwidth) can give model developers access to more GPUs per job to increase training throughput. For example, consider a model developer who seeks to maximize training throughput by using 32 A100 GPUs (c2 in Figure 1), but discovers that only 16 A100s (c0 in Figure 1) are currently available in one zone. Using heterogeneous GPU generations (e.g., c3 uses an additional 16 V100s in the zone) or multi-zone configurations (e.g., c4 uses 32 A100s spread across 2 zones within a region) increases throughput by 1.15$\times$ and 1.87$\times$, respectively, with a moderate increase in monetary cost per iteration (black line).
@@
However, supporting heterogeneous, geo-distributed resources introduces several challenges. First, heterogeneous GPU types and placements across zones exponentially expand the configuration search space. Searching for an optimal configuration requires jointly optimizing the resource allocation and job parallelization plan (e.g., data/pipeline/tensor parallelism degrees). For example, in Figure 1, although c5 uses the same number of GPUs as c3, it achieves much lower throughput due to a suboptimal parallelization plan. The search must also consider the cost caused by extra resources and data transfers. Cloud providers charge significant fees for inter-zone and inter-region data transfers (aws-interzone, ; gcp-interzone, ; az-interzone, ), which impacts geo-distributed configurations. In Figure 1, c6 uses the same number of GPUs and parallelization strategy as c4, but spreads training across regions (instead of zones within the same region), which increases cost.
Furthermore, as resource availability changes dynamically in datacenters, due to variable demand and node failures or preemptions, it is necessary to navigate this vast search space quickly  (Thorpe23Bamboo, ; wagenlander24tenplex, ). Figure 2 shows the varying number of A100 GPUs that we were able to allocate over an 8-hour period, out of the 8 A100s that we continuously requested in 2 different zones in Google Cloud. Such changes in resource availability require frequently re-evaluating the job configuration search space. Thus, model developers require a system to quickly navigate the large search space of heterogeneous (and homogeneous) configurations with dynamic resource availability, optimizing for the user’s performance-cost objective, while satisfying constraints (e.g., budget limits).
@@
Finally, after finding an appropriate resource allocation and parallelization plan for a training job given the available resources, model developers need to be able to run this job configuration in a distributed training framework (such as Megatron (shoeybi2020megatronlm, )). We find that optimal configurations for jobs running on heterogeneous resource topologies often include heterogeneous parallelism degrees per stage to load-balance the compute and memory capacity of different GPU types. Today’s state-of-the-art distributed training frameworks need to be adapted to support such heterogeneous job configurations. Furthermore, as resource availability can change frequently (e.g., when using spot instances (Thorpe23Bamboo, )), the training framework must be able to quickly reconfigure jobs.
@@
Existing systems do not adequately solve these challenges. First, current works do not co-optimize the resource allocation with the job parallelization plan. Instead, systems like Aceso (Liu24Aceso, ), Galvatron (Miao22Galvatron, ), and others in Table 1 expect the user to select a fixed resource allocation for which the system recommends a job parallelization plan. Most systems also do not consider heterogeneous resource topologies. Recent systems like Atlas (palak2024atlas, ), DTFM (yuan2023decentralized, ), Metis (Um24Metis, ), and FlashFlex (yan2024flashflex, ) optimize parallelization for heterogeneous GPUs or geo-distributed setups, but they suffer from prohibitively long search times (up to hours for configurations with 10s of GPUs) (Um24Metis, ), or suboptimal cost functions (yan2024flashflex, ; yuan2023decentralized, ), making them unsuitable for environments with dynamic resource availability. Second, existing systems rely on inaccurate simulators to estimate the training throughput and memory footprint of candidate configurations. For example, Varuna (Athlur22Varuna, ) overlooks significant memory sources (e.g., memory needed by the optimizer, communication, etc) when estimating memory footprint, hence recommending configurations that lead to OOM errors. Finally, state-of-the-art distributed training frameworks like Megatron-LM (shoeybi2020megatronlm, ) are slow to reconfigure jobs and do not support heterogeneous job parallelization plans or different microbatch sizes per GPU, which is necessary to maximize throughput and minimize cost in heterogeneous clusters.
@@
To this end, we propose Sailor, a system for efficient large-scale training over heterogeneous resources with dynamic availability. Sailor consists of three components: a configuration planner, a simulator, and a distributed training framework. The Sailor planner navigates the search space of resource allocations and job parallelization plan combinations. It recommends configurations that optimize a user-defined objective (e.g., max throughput or min cost) under constraints (e.g., max budget or min throughput). The planner considers heterogeneous GPU and machine types and geo-distributed setups. The planner uses the simulator to accurately model iteration time and memory footprint for any given configuration. Through a combination of dynamic programming and search space pruning with effective heuristics, the planner finds solutions within seconds even for allocations with 100s of GPUs and varying degrees of heterogeneity. This allows Sailor to quickly adapt plans based on resource availability. Finally, the Sailor training framework adds support for heterogeneous configurations to execute the planner’s configurations. It also improves fault-tolerance and elasticity by quickly adapting to changes in resource availability with efficient error detection and kill-free job reconfiguration. Together, these components enable Sailor to efficiently automate large-scale training in homogeneous, heterogeneous, and/or dynamic resource environments.
@@
We evaluate Sailor in various setups and compare it extensively to prior works. To the best of our knowledge, our work is the first to compare the major open-source ML training planners proposed to-date (Table 1) in homogeneous and heterogeneous scenarios. We show that Sailor can find resource allocations and job parallelization plans that result in higher throughput than baselines in both homogeneous and heterogeneous clusters. We show how Sailor can leverage heterogeneous resources to improve throughput by 1.1-2.87$\times$ compared to the heterogeneity-aware baselines (Metis, FlashFlex, AMP), while maintaining search times of 10s of seconds compared to minutes or hours needed by the baselines. We also show Sailor’s ability to increase performance and reduce cost in geo-distributed setups compared to DTFM by 5.9$\times$ and 9.8$\times$, respectively. Finally, we demonstrate Sailor’s ability to minimize monetary cost given throughput constraints, resulting in 40% cost savings compared to the second-best-performing baseline.
@@ 2.1. ML job parallelization strategies
Piper (Tarnawski21Piper, ) 3D, X, X, X ¡1 sec
AMP (li2022amp, ) 3D, X, X, X 14 sec
Varuna (Athlur22Varuna, ) 3D, X, X, X ¡ 1 sec
Oobleck (Jang2023Oobleck, ) 3D, X, X, X hours
Metis (Um24Metis, ) 3D, X, $\checkmark$, X hours
FlashFlex (yan2024flashflex, ) 3D, $\checkmark$, $\checkmark$, X 3 sec
Galvatron (Miao22Galvatron, ) 3D, X, X, X 10s of sec
Aceso (Liu24Aceso, ) 3D, X, X, X 200 sec
DTFM (yuan2023decentralized, ) 2D, $\checkmark$, X, $\checkmark$ 125 sec
Atlas (palak2024atlas, ) 3D, $\checkmark$, X, $\checkmark$ 100 sec
Sailor 3D, $\checkmark$,$\checkmark$,$\checkmark$ ¡ 1 sec
Table 1. Overview of distributed ML training planners. We omit planners that change ML training semantics (Wang23cocktailsgd, ). The Support column stands for: [degrees of parallelism supported, recommends resource allocation, supports heterogeneous GPU types, supports multi-zone]. Search time assumes a cluster of 128 A100 GPUs and OPT-350M model.
@@
Data Parallelism (DP): The model is replicated across workers, while the dataset is partitioned. At the end of each iteration, after each worker has computed the gradients on its own minibatch, the workers synchronize their gradients using an all-reduce collective (Sapio21innetwork, ).
Pipeline Parallelism (PP): The model is split into stages, with each stage consisting of a set of layers, and assigned to a worker or node. Workers operate at the granularity of a microbatch [^1] A minibatch is split in microbatches., performing forward and backward passes, sending activations to the next stage, and gradients to the previous stage. Due to inter-stage dependencies, pipeline parallelism is subject to bubbles, i.e., periods that a stage remains idle, waiting for others to complete (Huang19Gpipe, ). As a result, many approaches have been proposed to process multiple microbatches simultaneously and reduce bubbles (Athlur22Varuna, ; Narayanan19Pipedream, ).
Tensor Parallelism (TP): With tensor parallelism, a layer is divided across GPUs. After each GPU performs its local computations (both in forward and backward pass), the GPUs are synchronized using collectives such as all-reduce and all-gather. Since TP requires frequent communication, it requires very high interconnection bandwidths and is usually limited within a single node for reasonable throughput (shoeybi2020megatronlm, ).
@@ 2.2. Automating parallelization strategies
Determining the optimal degree of parallelism for each dimension (DP, PP, TP) is complex and greatly affects training throughput. Multiple systems, which we refer to as planners, automate this process (see Table 1). Given a fixed resource allocation (e.g., 16 nodes with 4 A100 GPUs each), model configuration (including hyperparameters like global batch size and learning rate), model profiling information (e.g., time for forward and backward pass for different configurations), and hardware characteristics (e.g., network bandwidth), planners explore different parallelization strategies, estimating the training time under different configurations, using some form of simulation. Planners apply techniques such as exhaustive search (Athlur22Varuna, ; Um24Metis, ), dynamic programming (Tarnawski21Piper, ; li2022amp, ), and integer linear programming (zheng2022alpa, ) to identify configurations that minimize training time.
@@ 3.1. Why use heterogeneous, geo-distributed GPUs?
Homogeneous high-end GPU clusters are scarce. The widespread adoption of ML, and hardware vendors’ inability to keep up with this pace, has significantly increased GPU demand. Several studies report limited GPU availability across public cloud providers (Zongheng23Skypilot, ; Cheng23availability, ; chugh23availability, ; strati24crossregion, ; Thorpe23Bamboo, ). For example, Figure 2 plots the availability of on-demand A100 GPUs in two zones in GCP. In one zone, it took 7 hours to allocate 8 A100 GPUs, while in the other zone, the requested number of GPUs was not attained within the 8-hour window. Our findings align with the AWS GPU availability plot from Guo et al. (guo2024cephalo, ), which shows that high-end GPUs such as A100 and H100 are difficult to aquire, while mid-tier GPUs (A10G, V100, T4) have higher but still limited and dynamic availability. Using GPUs across zones with heterogeneous types gives ML jobs the opportunity to use more GPUs to further increase throughput, as shown in Figure 1.
Power and cooling limits #GPUs per datacenter. From 2020 to 2025, the size of state-of-the-art ML models has increased by roughly 1200$\times$ (gpt-2, ; gpt-4, ; llama-4, ), while per-GPU memory capacity has only increased by 8$\times$ (nvidia-list, ). This requires hyperscalers to deploy more and more GPUs (ms-newAIdatacenter, ; meta-newAIdatacenter, ). However, the high power and cooling requirements of high-end GPUs limits the amount that can be deployed per datacenter (tapas, ). The next generation of ML models may need to train on GPUs across multiple availability zones (geodistributedblog1, ; geodistributedblog2, ; heatchallengeblog, ; palak2024improvingtrainingtimegpu, ).
Using old GPUs can reduce embodied carbon. Embodied carbon (greenhouse gas emissions associated with manufacturing to disposal) is a major source of datacenter emissions (greenSKU, ; carbonExplorer, ; metaEmbodiedCarbon, ). Although users prefer to allocate the latest GPUs for their ML jobs, older GPUs are abundant as the typical lifetime for ML servers in hyperscaler datacenters is $\sim$6 years (schneider2025lifecycleGoogleAI, ; greenSKU, ). Finding optimal ways to spread jobs across heterogeneous GPUs will enable leveraging older GPUs for longer to better amortize embodied carbon.

### v2

@@ Abstract.
Distributed Training
journalyear: 2025copyright: ccconference: ACM SIGOPS 31st Symposium on Operating Systems Principles; October 13–16, 2025; Seoul, Republic of Koreabooktitle: ACM SIGOPS 31st Symposium on Operating Systems Principles (SOSP ’25), October 13–16, 2025, Seoul, Republic of Koreadoi: 10.1145/3731569.3764839isbn: 979-8-4007-1870-0/2025/10ccs: Computing methodologies Machine learning
@@ 1. Introduction
GPUs are in high demand for large-scale Machine Learning (ML). As ML models continue to grow exponentially in size, they require an increasing number of GPUs to train and fine-tune. This high demand makes it difficult for model developers to allocate the desired number of accelerators to train models at high throughput in public clouds or enterprise clusters (Um et al., 2024; Strati et al., 2024a; Yang et al., 2023). Datacenters typically host a variety of GPU types and generations, spread across geographic regions (Cloud, 2024; Jia et al., 2022; Weng et al., 2022; Jeon et al., 2019; Wu et al., 2024). Yet model developers tend to restrict model training to homogeneous clusters of GPUs, since state-of-the-art distributed training frameworks like Megatron-LM (Shoeybi et al., 2020) and DeepSpeed (deepspeedai, 2025) assume homogeneous GPUs and inter-node bandwidth. The demand for large, homogeneous GPU clusters compounds the scarcity of high-end GPUs.
@@
Allowing a training job to run on heterogeneous GPU types and/or GPUs distributed across zones (i.e., with heterogeneous inter-node bandwidth) can give model developers access to more GPUs per job to increase training throughput. For example, consider a model developer who seeks to maximize training throughput by using 32 A100 GPUs (c2 in Figure 1), but discovers that only 16 A100s (c0 in Figure 1) are currently available in one zone. Using heterogeneous GPU generations (e.g., c3 uses an additional 16 V100s in the zone) or multi-zone configurations (e.g., c4 uses 32 A100s spread across 2 zones within a region) increases throughput by 1.15 and 1.87, respectively, with a moderate increase in monetary cost per iteration (black line).
However, supporting heterogeneous, geo-distributed resources introduces several challenges. First, heterogeneous GPU types and placements across zones exponentially expand the configuration search space. Searching for an optimal configuration requires jointly optimizing the resource allocation and job parallelization plan (e.g., data/pipeline/tensor parallelism degrees). For example, in Figure 1, although c5 uses the same number of GPUs as c3, it achieves much lower throughput due to a suboptimal parallelization plan. The search must also consider the cost caused by extra resources and data transfers. Cloud providers charge significant fees for inter-zone and inter-region data transfers (AWS, 2025; Cloud, 2025; Azure, 2025), which impacts geo-distributed configurations. In Figure 1, c6 uses the same number of GPUs and parallelization strategy as c4, but spreads training across regions (instead of zones within the same region), which increases cost.
Furthermore, as resource availability changes dynamically in datacenters, due to variable demand and node failures or preemptions, it is necessary to navigate this vast search space quickly  (Thorpe et al., 2023; Wagenländer et al., 2024). Figure 2 shows the varying number of A100 GPUs that we were able to allocate over an 8-hour period, out of the 8 A100s that we continuously requested in 2 different zones in Google Cloud. Such changes in resource availability require frequently re-evaluating the job configuration search space. Thus, model developers require a system to quickly navigate the large search space of heterogeneous (and homogeneous) configurations with dynamic resource availability, optimizing for the user’s performance-cost objective, while satisfying constraints (e.g., budget limits).
@@
Finally, after finding an appropriate resource allocation and parallelization plan for a training job given the available resources, model developers need to be able to run this job configuration in a distributed training framework (such as Megatron (Shoeybi et al., 2020)). We find that optimal configurations for jobs running on heterogeneous resource topologies often include heterogeneous parallelism degrees per stage to load-balance the compute and memory capacity of different GPU types. Today’s state-of-the-art distributed training frameworks need to be adapted to support such heterogeneous job configurations. Furthermore, as resource availability can change frequently (e.g., when using spot instances (Thorpe et al., 2023)), the training framework must be able to quickly reconfigure jobs.
@@
Existing systems do not adequately solve these challenges. First, current works do not co-optimize the resource allocation with the job parallelization plan. Instead, systems like Aceso (Liu et al., 2024), Galvatron (Miao et al., 2022), and others in Table 1 expect the user to select a fixed resource allocation for which the system recommends a job parallelization plan. Most systems also do not consider heterogeneous resource topologies. Recent systems like Atlas (Palak et al., 2024a), DTFM (Yuan et al., 2023), Metis (Um et al., 2024), and FlashFlex (Yan et al., 2024) optimize parallelization for heterogeneous GPUs or geo-distributed setups, but they suffer from prohibitively long search times (up to hours for configurations with 10s of GPUs) (Um et al., 2024), or suboptimal cost functions (Yan et al., 2024; Yuan et al., 2023), making them unsuitable for environments with dynamic resource availability. Second, existing systems rely on inaccurate simulators to estimate the training throughput and memory footprint of candidate configurations. For example, Varuna (Athlur et al., 2022) overlooks significant memory sources (e.g., memory needed by the optimizer, communication, etc) when estimating memory footprint, hence recommending configurations that lead to OOM errors. Finally, state-of-the-art distributed training frameworks like Megatron-LM (Shoeybi et al., 2020) are slow to reconfigure jobs and do not support heterogeneous job parallelization plans or different microbatch sizes per GPU, which is necessary to maximize throughput and minimize cost in heterogeneous clusters.
@@
To this end, we propose Sailor, a system for efficient large-scale training over heterogeneous resources with dynamic availability. Sailor  [^1] Sailor is available at https://github.com/eth-easl/sailor consists of three components: a configuration planner, a simulator, and a distributed training framework. The Sailor planner navigates the search space of resource allocations and job parallelization plan combinations. It recommends configurations that optimize a user-defined objective (e.g., max throughput or min cost) under constraints (e.g., max budget or min throughput). The planner considers heterogeneous GPU and machine types and geo-distributed setups. The planner uses the simulator to accurately model iteration time and memory footprint for any given configuration. Through a combination of dynamic programming and search space pruning with effective heuristics, the planner finds solutions within seconds even for allocations with 100s of GPUs and varying degrees of heterogeneity. This allows Sailor to quickly adapt plans based on resource availability. Finally, the Sailor training framework adds support for heterogeneous configurations to execute the planner’s configurations. It also adds support for fault tolerance and elasticity, enabling adaptation to changes in resource availability. Together, these components enable Sailor to efficiently automate large-scale training in homogeneous, heterogeneous, and/or dynamic resource environments.
@@
We evaluate Sailor in various setups and compare it extensively to prior works. To the best of our knowledge, our work is the first to compare the major open-source ML training planners proposed to-date (Table 1) in homogeneous and heterogeneous scenarios. We show that Sailor can find resource allocations and job parallelization plans that result in higher throughput than baselines in both homogeneous and heterogeneous clusters. We show how Sailor can leverage heterogeneous resources to improve throughput by 1.1-2.87 compared to the heterogeneity-aware baselines (Metis, FlashFlex, AMP), while maintaining search times of 10s of seconds compared to minutes or hours needed by the baselines. We also show Sailor’s ability to increase performance and reduce cost in geo-distributed setups compared to DTFM by 5.9 and 9.8, respectively. Finally, we demonstrate Sailor’s ability to minimize monetary cost given throughput constraints, resulting in 40% cost savings compared to the second-best-performing baseline.
@@ 2.1. ML job parallelization strategies
Piper (Tarnawski et al., 2021) 3D, X, X, X ¡1 sec
AMP (Li et al., 2022) 3D, X, , X 14 sec
Varuna (Athlur et al., 2022) 3D, X, X, X ¡ 1 sec
Oobleck (Jang et al., 2023) 3D, X, X, X hours
Metis (Um et al., 2024) 3D, X, , X hours
FlashFlex (Yan et al., 2024) 3D, , , X 3 sec
Galvatron (Miao et al., 2022) 3D, X, X, X 10s of sec
Aceso (Liu et al., 2024) 3D, X, X, X 200 sec
DTFM (Yuan et al., 2023) 2D, , X, 125 sec
Atlas (Palak et al., 2024a) 3D, , X, 100 sec
Sailor 3D, ,, ¡ 1 sec
Table 1. Overview of distributed ML training planners. We omit planners that change ML training semantics (Wang et al., 2023). The Support column stands for: [degrees of parallelism supported, recommends resource allocation, supports heterogeneous GPU types, supports multi-zone]. Search time assumes a cluster of 128 A100 GPUs and OPT-350M model.
@@
Data Parallelism (DP): The model is replicated across workers, while the dataset is partitioned. At the end of each iteration, the workers synchronize their gradients using an all-reduce collective (Sapio et al., 2021).
Pipeline Parallelism (PP): The model is split into stages, with each stage consisting of a set of layers, and assigned to a worker or node. Workers operate at the granularity of a microbatch [^2] A minibatch is split in microbatches., performing forward and backward passes, sending activations to the next stage, and gradients to the previous stage. Due to inter-stage dependencies, pipeline parallelism is subject to bubbles, i.e., periods that a stage remains idle, waiting for others to complete (Huang et al., 2019). As a result, many approaches have been proposed to process multiple microbatches simultaneously and reduce bubbles (Athlur et al., 2022; Narayanan et al., 2019).
Tensor Parallelism (TP): With tensor parallelism, a layer is divided across GPUs. After each GPU performs its local computations (both in forward and backward pass), the GPUs are synchronized using collectives such as all-reduce and all-gather. Since TP requires frequent communication, it requires very high interconnection bandwidths and is usually limited within a single node for reasonable throughput (Shoeybi et al., 2020).
@@ 2.2. Automating parallelization strategies
Determining the optimal degree of parallelism for each dimension (DP, PP, TP) is complex and greatly affects training throughput. Multiple systems, which we refer to as planners, automate this process (see Table 1). Given a fixed resource allocation (e.g., 16 nodes with 4 A100 GPUs each), model configuration (including hyperparameters like global batch size and learning rate), model profiling information (e.g., time for forward and backward pass for different configurations), and hardware characteristics (e.g., network bandwidth), planners explore different parallelization strategies, estimating the training time under different configurations, using some form of simulation. Planners apply techniques such as exhaustive search (Athlur et al., 2022; Um et al., 2024), dynamic programming (Tarnawski et al., 2021; Li et al., 2022), and integer linear programming (Zheng et al., 2022) to identify configurations that minimize training time.
@@ 3.1. Why use heterogeneous, geo-distributed GPUs?
Homogeneous high-end GPU clusters are scarce. The widespread adoption of ML, and hardware vendors’ inability to keep up with this pace, has significantly increased GPU demand. Several studies report limited GPU availability across public cloud providers (Yang et al., 2023; Cheng et al., 2023; Chugh et al., 2023; Strati et al., 2024a; Thorpe et al., 2023). For example, Figure 2 plots the availability of on-demand A100 GPUs in two zones in GCP. In one zone, it took 7 hours to allocate 8 A100 GPUs, while in the other zone, the requested number of GPUs was not attained within the 8-hour window. Our findings align with the AWS GPU availability plot from Guo et al. (Guo et al., 2024), which shows that high-end GPUs such as A100 and H100 are difficult to aquire, while mid-tier GPUs (A10G, V100, T4) have higher but still limited and dynamic availability. Using GPUs across zones with heterogeneous types gives ML jobs the opportunity to use more GPUs to further increase throughput, as shown in Figure 1.
Power and cooling limits #GPUs per datacenter. From 2020 to 2025, the size of state-of-the-art ML models has increased by roughly 1200 (Wikipedia, 2025a, b; Meta, 2025), while per-GPU memory capacity has only increased by 8 (Wikipedia, 2025c). This requires hyperscalers to deploy more and more GPUs (tom’s Hardware, 2025; Hardware, 2025). However, the high power and cooling requirements of high-end GPUs limits the amount that can be deployed per datacenter (Stojkovic et al., 2025). The next generation of ML models may need to train on GPUs across multiple availability zones (Semaphor, 2024; Baseline, 2024; Linkedin, 2024; Palak et al., 2024b).
Using old GPUs can reduce embodied carbon. Embodied carbon (greenhouse gas emissions associated with manufacturing to disposal) is a major source of datacenter emissions (Wang et al., 2024; Acun et al., 2023; Rivalin et al., 2025). Although users prefer to allocate the latest GPUs for their ML jobs, older GPUs are abundant as the typical lifetime for ML servers in hyperscaler datacenters is 6 years (Schneider et al., 2025; Wang et al., 2024). Finding optimal ways to spread jobs across heterogeneous GPUs will enable leveraging older GPUs for longer to better amortize embodied carbon.